# Improving Detection in Aerial Images by Capturing Inter-Object Relationships

### v1

@@ Abstract
In many image domains, the spatial distribution of objects in a scene exhibits meaningful patterns governed by their semantic relationships. In most modern detection pipelines, however, the detection proposals are processed independently, overlooking the underlying relationships between objects. In this work, we introduce a transformer-based approach to capture these inter-object relationships to refine classification and regression outcomes for detected objects. Building on two-stage detectors, we tokenize the region of interest (RoI) proposals to be processed by a transformer encoder. Specific spatial and geometric relations are incorporated into the attention weights and adaptively modulated and regularized. Experimental results demonstrate that the proposed method achieves consistent performance improvement on three benchmarks including DOTA-v1.0, DOTA-v1.5, and HRSC 2016, especially ranking first on both DOTA-v1.5 and HRSC 2016. Specifically, our new method has an increase of 1.59 mAP on DOTA-v1.0, 4.88 mAP on DOTA-v1.5, and 2.1 mAP on HRSC 2016, respectively, compared to the baselines.
Keywords:
Aerial image processing Object detection Inter-object relationship utilization
@@ 1 Introduction
(a) Baseline
(b) Our method
Figure 1: Visualization of a motivating example. (a) Detections obtained by the baseline [10] have erroneous identifications: the upper left image shows a false detection of ship on top of an airplane; the bottom left image shows an incorrect airplane detection with unrealistic size. (b) Improved results obtained by our method. The false positives are effectively addressed. Highlighting the importance of considering inter-object relationship in detection.
@@
Nevertheless, despite these advancements, most present-day detection models classify and localize objects independently [10, 27, 32], possibly due to the lack of an effective tool for modeling the co-presence of an arbitrary number of objects in an image. In other words, the spatial and semantic relationships among objects are not fully captured, often leading to false detections that overlook contextual dependencies and inter-object dynamics. As a motivating example, Fig. 1 illustrates the challenge of detecting each object instance based solely on its features, without considering these critical relationships. Aerial images, in particular, offer a unique setting where objects generally share the same plane, with little occlusion and perspective distortion, and therefore have stable inter-object relationships.
Addressing the aforementioned gap, in this paper, we propose a Transformer-based model on top of two-stage detectors to effectively capture and leverage the inter-object relationships. Concretely, we organize the Region of Interest (RoI) [8, 23] feature maps proposed in the first stage and the independent detection results on them into embeddings. The embeddings are then fed into a transformer where the features of candidate detections interact and aggregate. However, the self-attention module in ordinary Transformers, which computes the pairwise attention weights as dot products of embeddings, does not capture the spatial and geometric relationship directly. To overcome this, we design and incorporate additional encodings and attention functions, weighing the mutual influence between objects according to distances. The attention functions are adaptive to the scales and densities of the object distribution, which is crucial for the model to generalize across different image scenarios. We conduct comprehensive experiments on DOTA-v1.0, DOTA-v1.5 [26], and HRSC2016 [18] to evaluate the efficacy of our model, achieving an improvement of 1.59, 4.88, and 2.1 mAP over the baseline.
@@ 2 Related Works
2.1 Object Detection in Aerial Images
@@ 3 Methodology
We build our method on the two-stage object detection framework presented in [10]. We start by following the original pipeline to obtain features and preliminary detections, which are then transformed into embeddings and input into the Transformer with additional encodings. To better leverage the Transformer, we introduce a novel attention function on top of the common scaled dot product and a set of spatial relations. It aims to reflect the degree of correlation between objects based on distances in the image, emphasizing neighboring detections while being aware of object scale and density. Eventually, we perform another detection on the features given by the Transformer to obtain the final results. The overview of our model is shown in Fig. 2.
Figure 2: Overview of our model. Left: the detection pipeline where the Transformer Encoder is trained to capture the relationships between RoI Tokens. Right: calculation of the adaptive attention weights. Slashes indicate gradient stopping.
@@ 3.2 Spatial and Geometric Relations
Name Formula Description
@@ 3.3 Adaptive Attention Weights
3.3.1 Spatial Distance
The scale of aerial images could be several kilometers. To focus more on the neighboring objects that are expected to be more relevant, we introduce a distance-decaying coefficient for each object pair:
$A_{ij}=\underbrace{\exp(-(\epsilon_{i}d_{ij})^{2}/\sigma^{2})}_{\text{distance, scale and density}}\circ\ \underbrace{\mathbf{1}\{\text{IoU}_{ij}<\delta\}}_{\text{overlapping}}$ (3)
where $d_{ij}=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$ is the pair-wise distance, $\circ$ the element-wise product, and $\mathbf{1}\{\cdot\}$ the indicator function. $\epsilon_{i}$ is detailed next. $\sigma$ is a hyperparameter.
3.3.2 Object Scale and Density
It is a natural intuition that the influence of one object on another relates to the scale (size) of the object, e.g., smaller objects tend to be more influenced by nearby objects, whereas larger objects need to capture the impact at longer ranges. The density around a proposed detection is another important factor. We assume that when there are fewer other RoIs around a detection (i.e., lower density), it should capture the influence of RoIs from further away. In contrast, in areas with high density (many RoIs), it should mainly interact with nearby RoIs. To qualitatively model these factors, we compute $\epsilon_{i}$ as:
$\epsilon_{i}=\frac{S}{\sqrt{w_{i}h_{i}}}\times\exp(\bar{\rho}_{i})$ (4)
where $S$ is a global (dataset-wide) scale factor determined by the input image.
@@ 3.3.2 Object Scale and Density
Figure 3: Illustration of the adaptive attention weights w.r.t one RoI. Areas with high opacity indicate higher weights $\epsilon_{i}$. In the second row, the left image demonstrates a tennis court with a larger area and fewer surrounding objects, resulting in a smaller $\epsilon_{i}$. This leads to a slower decay in Eq. 3 as distance increases, enabling the model to capture relationships with more distant objects. Conversely, the right image depicts a car with a smaller area and a higher density of surrounding objects, which entails a larger $\epsilon_{i}$, resulting in a faster decay of Eq. 3 and thus a focus on objects in closer proximity.
@@ 4.1 Dataset
DOTA is a large-scale aerial object detection dataset.
@@
DOTA-v1.5 uses the same image set but with increased annotations. This version features 402,089 instances and introduces an additional category, Container Crane (CC), broadening the dataset’s applicability in aerial object detection.
@@ 4.2 Implementation Details
Our implementation is based on the MMRotate [36] library and adopts ReDet’s framework and hyperparameter settings. We train our model for $12$ epochs using the AdamW [19] optimizer with an initial learning rate of $41e-$ 4, reduced to $51e-$ 5 and $61e-$ 6 at epochs 8 and 11. We also use a weight decay of 0.05. The experiments were conducted using two RTX 3090 GPUs.
The Transformer module consists of 6 encoder layers, similar to the ViT structure, and integrates sinusoidal two-dimensional absolute position encoding, hyperparameters $\sigma$ is set to 4. A dropout rate of 0.1 is employed during the training phase of the Transformer.
Table 2: Results of each object class on the DOTA-v1.5 dataset. We highlight the best in blue.
Method PL BD BR GTF SV LV SH TC BC ST SBF RA HA SP HC CC mAP RetinaNet-O[16] 71.43 77.64 42.12 64.65 44.53 56.79 73.31 90.84 76.02 59.96 46.95 69.24 59.65 64.52 48.06 0.83 59.16 RF. R-CNN [23] 72.20 76.43 47.58 69.91 51.99 70.52 80.27 90.87 79.16 68.63 59.57 72.34 66.44 66.07 55.29 6.87 64.63 Mask R-CNN[11] 76.84 73.51 49.90 57.80 51.31 71.34 79.75 90.46 74.21 66.07 46.21 70.61 63.07 64.46 57.81 9.42 62.67 HTC [3] 77.80 73.67 51.40 63.99 51.54 73.31 80.31 90.48 75.21 67.34 48.51 70.63 64.84 64.48 55.87 5.15 63.40 RoI-Trans. [6] 72.27 81.95 54.47 70.02 52.49 76.31 81.03 90.90 84.19 69.12 62.85 72.73 68.67 65.89 57.09 7.12 66.69 ReDet [10] 79.20 82.81 51.92 71.41 52.38 75.73 80.92 90.83 75.81 68.64 49.29 72.03 73.36 70.55 63.33 11.53 66.86 Ours 80.79 85.58 53.80 71.74 52.72 77.84 88.71 90.89 86.24 74.73 64.97 73.31 76.76 72.99 73.84 22.87 71.74
Table 3: More experiments on DOTA-v1.5. We highlight our method’s improvements in blue
. Method mAP RoI Transformer 66.69 RoI Transformer + Ours 68.61(+1.92) Oriented RCNN 65.94 Oriented RCNN + Ours 66.38(+0.44) Rotated Faster RCNN 64.63 Rotated Faster RCNN + Ours 65.28(+0.65) ReDet 66.86 ReDet + Ours 71.74(+4.88)
@@ 4.3 Comparison with Baselines
First, we evaluate our model against the baselines on DOTA-v1.0, DOTA-v1.5, and HRSC2016 to demonstrate the efficacy of the proposed method. The results are shown in Tab. 2, Tab. 4, respectively. To further illustrate the versatility of our approach, we conducted additional experiments on DOTA-v1.5 with our method applied to various models, yielding consistent improvements, as shown in Tab. 3. These results demonstrate that our method consistently outperforms the baselines across different datasets. Notably, however, the improvement achieved in HRSC2016 is marginal compared to that on DOTA-v1.5. This is possibly due to the number of instances in a single image being much fewer in HRSC (typically less than 4), thus there are limited opportunities to leverage the inter-object relationships. These findings suggest that our model’s strengths are most pronounced in scenarios rich in object interactions and contextual dynamics, aligning with our design’s focus on capturing and utilizing inter-object relationships.
Table 4: Results in COCO style on DOTA-v1.0 and HRSC2016. We highlight our method’s improvements in blue
. Dataset Method AP50 AP75 mAP DOTA-v1.0 ReDet 76.25 50.86 47.11 Ours 77.84(+1.59) 51.42(+0.56) 48.32(+1.21) HRSC2016 ReDet 90.46 89.46 70.41 Ours 90.49(+0.03) 89.67(+0.21) 72.51(+2.10)
Table 5: Detection performance with and without preliminary detection training.
Table 6: Effect of the components in computing the attention weights via Eq. 7
pre cls supervision mAP
w/o 69.86

### v2

# Context-Aware Aerial Object Detection: Leveraging Inter-Object and Background Relationships

@@ Abstract
In most modern object detection pipelines, the detection proposals are processed independently given the feature map. Therefore, they overlook the underlying relationships between objects and the surrounding background, which could have provided additional context for accurate detection. Because aerial imagery is almost orthographic, the spatial relations in image space closely align with those in the physical world, and inter-object and object-background relationships become particularly significant. To address this oversight, we propose a framework that leverages the strengths of Transformer-based models and Contrastive Language-Image Pre-training (CLIP) features to capture such relationships. Specifically, Building on two-stage detectors, we treat Region of Interest (RoI) proposals as tokens, accompanied by CLIP Tokens obtained from multi-level image segments. These tokens are then passed through a Transformer encoder, where specific spatial and geometric relations are incorporated into the attention weights, which are adaptively modulated and regularized. Additionally, we introduce self-supervised constraints on CLIP Tokens to ensure consistency. Extensive experiments on three benchmark datasets demonstrate that our approach achieves consistent improvements, setting new state-of-the-art results with increases of 1.37 mAP50 on DOTA-v1.0, 5.30 mAP50 on DOTA-v1.5, 2.30 mAP50 on DOTA-v2.0 and 3.23 mAP50 on DIOR-R.
@@ 1 Introduction
(a) ReDet
(b) ReDet + Ours
Figure 1: Visualization of a motivating example. (a) Detections obtained by the ReDet [11] have erroneous identifications: the upper left image shows a false detection of ship on top of an airplane; the bottom left image shows an incorrect airplane detection with unrealistic size. (b) Improved results obtained by our method. The false positives are effectively addressed. Highlighting the importance of considering inter-object relationship and background context in detection.
@@
Nevertheless, despite these advancements, most present-day detection models classify and localize objects independently [11, 34, 38], possibly due to the lack of an effective tool for modeling the co-presence of an arbitrary number of objects in an image. In other words, the spatial and semantic relationships among objects are not fully captured, often leading to false detections that overlook surrounding contextual dependencies and inter-object dynamics. As a motivating example, Fig. 1 illustrates the challenge of detecting each object instance based solely on its features, without considering these critical relationships. Aerial images, in particular, offer a unique setting where objects generally share the same plane, with little occlusion and perspective distortion, and therefore have stable inter-object and surrounding context relationships. Meanwhile, we posit that knowledge of an object’s background context can provide useful information and therefore significantly improve detection. For instance, an area that appears as a green field might be presumed a playground; however, if adjacent to an airport, such an assumption would be reconsidered. Unfortunately, most datasets lack annotations for background information. The semantics of the background are complex and difficult to annotate due to the highly irregular spatial distribution.
In this paper, we propose a Transformer-based model on top of two-stage detectors to effectively capture and leverage the inter-object relationships and semantic background context. Concretely, we organize the Region of Interest (RoI) [9, 26] feature maps proposed in the first stage and the independent detection results on them into embeddings. The embeddings are then fed into a transformer where the features of candidate detections interact and aggregate. However, the self-attention module in ordinary Transformers, which computes the pairwise attention weights as dot products of embeddings, does not capture the spatial and geometric relationship directly. To overcome this, we design and incorporate additional encodings and attention functions, weighing the mutual influence between objects according to distances. The attention functions are adaptive to the scales and densities of the object distribution, which is crucial for the model to generalize across different image scenarios.
To further incorporate object-background relationships, we leverage CLIP [24], a powerful multimodal model renowned for its cross-modal understanding capabilities, to integrate background information into detectors. Utilizing the image and text encoders of a pre-trained CLIP model, we divide the image into patches to be queried by pre-specified descriptions and then cast them as tokens alongside the RoI tokens.
Aerial images offer complex scenes with numerous objects on a single plane, where spatial and inter-object relationships are more explicit. The richer background context in such scenarios further supports the strengths of our approach. We validate the effectiveness of our method through comprehensive experiments on DOTA-v1.0, DOTA-v1.5, DOTA-v2.0[33], and DIOR-R [6], achieving an improvement of 1.37, 5.30, 2.30 and 3.23 mAP50 over the baseline.
@@
  • We propose to use CLIP for integrating background context into the detection pipeline. We introduce multi-scale, hierarchical CLIP patches, generating CLIP Tokens and facilitating the flow of semantic information across different levels, thereby improving information fusion.
@@ 2 Related Works
2.1 Aerial Object Detection
@@ 2.2 Capturing Inter-Object Relationships
2.3 CLIP Features
The CLIP (Contrastive Language-Image Pretraining) [24] model is trained on a large corpus of image-text pairs and could be used to extract semantic information from images. CLIP has promoted the field of computer vision and is widely applicated in traditional vision tasks[24]. For example, the model and concepts of CLIP have been applied to a variety of other visual tasks, such as object detection, video action recognition and scene graph generation, showcasing its broad applicability and adaptability[32, 44, 30, 29]. Its zero-shot classification capability allows it to accurately categorize images without additional fine-tuning on specific datasets. This feature is particularly useful in tasks and fields where labeled data is scarce. Building upon the CLIP model, RegionCLIP[45] focuses on specific image regions for detailed semantic analysis. By focusing on distinct image regions, RegionCLIP can provide more precise and contextually relevant semantic information, which is critical for tasks that require the understanding of spatial relationships and localized features. In addition to CLIP and RegionCLIP, uniDetector[31] shows the advancement of leveraging the rich information from both visual and language modalities. As an object detection model, uniDetector combines the global context of transformers with the local feature extraction of CNNs, which is beneficial for visual tasks that need to handle objects of varying orientations and scales. Besides, the rich information from both visual and language modalities endows uniDetector with the ability to recognize open-vocabulary objects.
@@ 3 Methodology
We build our method on the two-stage object detection framework presented in [11]. We start by following the original pipeline to obtain features and preliminary detections, which are then transformed into RoI tokens. Then we segment images into multi-scale patches and use CLIP to generate multi-level CLIP features, and then transform them into CLIP tokens. These tokens are input into the Transformer with additional encodings. To better leverage the Transformer, we introduce a novel attention function on top of the common scaled dot product and a set of spatial relations. It aims to reflect the degree of correlation between objects based on distances in the image, emphasizing neighboring detections while being aware of object scale and density. Moreover, we introduce self-supervised constraints on CLIP Tokens, providing additional supervised signals. Eventually, we perform another detection on the features given by the Transformer to obtain the final results. The overview of our model is shown in Fig. 2.
Figure 2: Overview of our model. The model utilizes a two-stage detection framework where features are converted into RoI tokens. Multi-scale patches generate CLIP tokens to capture background context, as shown in (a). Both tokens are processed by a Transformer encoder (b) with spatially aware attention, enhancing inter-object relationships based on distance and scale. Self-supervised constraints on CLIP tokens aid background classification, leading to refined detections with supervised signals $L_{\text{reg}}+L_{\text{cls}}$ and $L_{\text{self}}$.
@@ 3.1 RoI Tokens
3.2 CLIP Tokens
Besides RoI tokens, we additionally introduce CLIP Tokens to capture and articulate the multi-scale semantic context offered by the background. We use a CLIP model fine-tuned on the RSCID dataset fine-tuned on the RSCID dataset [21]. We divide each image into patches of size $1$, $\frac{1}{2}$, $\frac{1}{4}$, and $\frac{1}{8}$, with strides 0, $\frac{1}{4}$, $\frac{1}{4}$, $\frac{1}{8}$, respectively. This results in 1, $5\times 5$, $7\times 7$, and $8\times 8$ patches. The patches are resized and passed through the CLIP image encoder to get their image embeddings $\mathbf{f}_{\text{image}}\in\mathrm{R}^{139\times 512}$. We also compute text embeddings using a set of pre-defined descriptions of the format "Aerial photograph of [object]", where object is selected from a range of natural and human landscapes such as forest, ocean, farmland, road, and airport. This results in $\mathbf{f}_{\text{text}}\in\mathrm{R}^{36\times 512}$ from 36 descriptions. Incorporating the semantic information from text embeddings, CLIP tokens analogous to Eq. 1 with $\mathbf{c}=\mathbf{f}_{\text{image}}\cdot\mathbf{f}_{\text{text}}^{T}$ and $\mathbf{w}$, $\mathbf{h}$ defined similarly according to patch sizes and locations. Moreover, to better combine information from patches of different sizes, we fuse them in a way similar to FPN. This part is depicted in Fig. 2 (a).
@@ 3.3 Spatial and Geometric Relations
Symbol Formula Description
@@ 3.4 Adaptive Attention Weights
Spatial Distance, Scale, and Density.
It is a natural intuition that the influence of one object on another relates to the distance between them and the relative scales (sizes) of the object. For example, closer objects are assumed to have a stronger correlation than distant pairs, and smaller objects tend to be more influenced by nearby objects, whereas larger objects need to capture the impact at longer ranges. The density around a proposed detection is another important factor. We assume that when there are fewer other RoIs around a detection (i.e., lower density), it should capture the influence of RoIs from further away. To qualitatively model these factors, we compute $\epsilon_{i}$ as:
$A_{ij}=$ $\underbrace{\exp(-(\epsilon_{i}d_{ij})^{2}/\sigma^{2})}_{\text{distance, scale and density}}\circ\ \underbrace{\mathbf{1}\{\text{IoU}_{ij}<\delta\}}_{\text{overlapping}},$ (3)
$\epsilon_{i}=$ $\frac{S}{\sqrt{w_{i}h_{i}}}\times\exp(\bar{\rho}_{i}),$
where $d_{ij}=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$ is the pair-wise distance, $\circ$ the element-wise product, and $\mathbf{1}\{\cdot\}$ the indicator function. $\epsilon_{i}$ is detailed next. $\sigma$ is a hyperparameter, and $S$ is a global (dataset-wide) scale factor determined by the input image.
@@
3.5 Loss Function
We utilized a preliminary stage classification loss $L_{cls}^{pre}$, and final stage detection losses $L_{cls}$ and $L_{reg}$. Furthermore, to constrain CLIP tokens, we employed a self-supervised loss. The overall loss function is given by:
$L=L_{cls}+L_{reg}+\gamma L_{cls}^{pre}+\lambda L_{self}$ (7)
Where $L_{reg}$ is the standard bounding box regression loss, and $L_{cls}$ is the cross-entropy loss used in both stages. $L_{self}$ is the MSE loss between the background classification output $c_{clip}$ and its ground truth $c_{clip}^{GT}$. $\gamma$ and $\lambda$ are hyperparameters.
@@ 4.1 Dataset
Method PL BD BR GTF SV LV SH TC BC ST SBF RA HA SP HC CC mAP50
RetinaNet-O[18] 71.43 77.64 42.12 64.65 44.53 56.79 73.31 90.84 76.02 59.96 46.95 69.24 59.65 64.52 48.06 0.83 59.16
RF. R-CNN [26] 72.20 76.43 47.58 69.91 51.99 70.52 80.27 90.87 79.16 68.63 59.57 72.34 66.44 66.07 55.29 6.87 64.63
Mask R-CNN[12] 76.84 73.51 49.90 57.80 51.31 71.34 79.75 90.46 74.21 66.07 46.21 70.61 63.07 64.46 57.81 9.42 62.67
HTC [3] 77.80 73.67 51.40 63.99 51.54 73.31 80.31 90.48 75.21 67.34 48.51 70.63 64.84 64.48 55.87 5.15 63.40
RoI-Trans. [7] 72.27 81.95 54.47 70.02 52.49 76.31 81.03 90.90 84.19 69.12 62.85 72.73 68.67 65.89 57.09 7.12 66.69
ReDet [11] 79.20 82.81 51.92 71.41 52.38 75.73 80.92 90.83 75.81 68.64 49.29 72.03 73.36 70.55 63.33 11.53 66.86
FRED [17] 79.60 81.44 52.60 72.57 58.07 74.82 86.12 90.81 82.13 74.84 53.37 72.93 69.51 69.91 54.82 19.27 68.30
DCFL [35] - - - - 57.31 - 86.60 - - 76.55 - - - - - - 70.24
Ours (ReDet) 80.17 83.71 54.28 70.31 52.80 77.42 88.46 90.84 86.02 75.04 68.19 73.09 76.94 74.51 73.86 28.86 72.16
Table 2: Results of each object class on the DOTA-v1.5 dataset. We use the standard 1x training schedule for fair comparisons.
Dataset Method mAP50 mAP75 mAP50:95
DOTA-v1.0 ReDet 76.25 50.86 47.11
Ours 77.62 (+1.37) 52.18 (+1.32) 48.67 (+1.56)
HRSC2016 ReDet 90.46 89.46 70.41
Ours 90.49 (+0.03) 89.67 (+0.21) 72.51 (+2.10)
Table 3: Results in COCO style on DOTA-v1.0 and HRSC2016.
Method DOTA-v1.0 DOTA-v1.5 DOTA-v2.0 DIOR-R
Baseline 76.25 66.86 53.28 65.79
Ours 77.62 (+1.37) 72.16 (+5.30) 55.58 (+2.30) 69.02 (+3.23)
Table 4: mAP50 results on DOTA-v1.0, DOTA-v1.5, DOTA-v2.0, and DIOR-R. For DOTA-v2.0, we use Oriented R-CNN as the baseline due to out-of-memory issues with the original ReDet. For the other datasets, ReDet is used as the baseline.
@@
DOTA-v1.5 uses the same image set but with increased annotations. This version features 402,089 instances and introduces an additional category, Container Crane (CC), broadening the dataset’s applicability. It also includes annotations for a greater number of small objects, some of which have areas smaller than 10 pixels, further enhancing the dataset’s complexity.
DOTA-v2.0 further expands the datasets to 11,268 images and 1,793,658 instances, with two additional categories, Airport (AP) and Helipad (HP).
DIOR-R is a refined version of the original DIOR dataset, specifically re-annotated with rotated bounding boxes to enhance the detection of object orientation and shape in aerial images. It consists of 23,463 high-resolution images and 190,288 annotated instances, covering 20 diverse object categories, including vehicles, airplanes, ships, and more.
@@ 4.2 Implementation Details
Our implementation is based on the MMRotate [46] library and adopts ReDet’s framework and hyperparameter settings. We train our model for $12$ epochs using the AdamW [20] optimizer with an initial learning rate of $1e^{-4}$, reduced to $1e^{-5}$ and $1e^{-6}$ at epochs 8 and 11. We also use a weight decay of 0.05. The experiments were conducted using two RTX 3090 GPUs.
The Transformer module consists of 6 encoder layers, similar to the ViT structure, and integrates sinusoidal two-dimensional absolute position encoding, hyperparameters $\sigma$ is set to 4. A dropout rate of 0.1 is employed during the training phase of the Transformer. In the loss function, we set $\gamma$ as 1, and $\lambda$ as 10.
@@ 4.3 Comparison with Baselines
First, we evaluate our model against the baselines on DOTA-v1.0, DOTA-v1.5, DOTA-v2.0, DIOR-R and HRSC2016 to demonstrate the efficacy of the proposed method. The results are shown in Tab. 3, and Tab. 4, respectively. These results demonstrate that our method consistently outperforms the baselines across different datasets. Notably, however, the improvement achieved in HRSC2016 is marginal compared to that on DOTA-v1.5 and DOTA-v2.0. This is possibly due to the number of instances in a single image being much fewer in HRSC (typically less than 4), thus there are limited opportunities to leverage the inter-object relationships. These findings suggest that our model’s strengths are most pronounced in scenarios rich in object interactions and contextual dynamics, aligning with our design’s focus on capturing and utilizing inter-object relationships.
4.4 Ablation Study
Preliminary Detection Phase.
Compared to the standard detection pipeline, our model incorporates two detection heads - placed before and after the Transformer module. The output from the initial detection phase, dubbed preliminary detection, includes a classification result (parameterized as a softmax distribution) from the first head, which forms a component of the RoI token. We posit that knowing the class information with uncertainties would help with reasoning about the inter-object relationships. To empirically validate this hypothesis, we compared the performance of our model with and without training the first detection head. As Tab. 5 shows, although solely incorporating the Transformer offers an improvement of mAP50 to the baseline, omitting the preliminary detection leads to a notable decline in performance. This suggests that relying only on the Transformer for RoIs to interact lacks efficacy. In contrast, the explicit inclusion of preliminary classification data, despite its potential inaccuracies, enhances the model’s ability to reason about semantical and contextual relationships. The results underscore the value of early classification cues in guiding the relational reasoning process within our proposed architecture.
Spatial-Geometric Relations and Adaptive Attention Weights.
The different terms presented in Tab. 1 characterize various aspects of the spatial and geometric relationships among objects (RoI Tokens) within an image. In this section, we aim to empirically evaluate the individual and collective contributions of these spatial and geometric relational terms to the overall performance of our detection model. As shown in Tab. 7, IoU and rel. area contribute the most. Intuitively, they are particularly helpful when reasoning about the co-occurrence and spatial arrangement of objects. For example, IoU helps to disambiguate the overlapping, potentially duplicate or conflicting detections. Similarly, relative area aids in discerning the size relationship between objects. Consequently, our method can effectively solve the problem in the motivation example. See Sec. 5.1 for details.
As mentioned in Sec. 3.4, making the attention weights adaptive to specific RoI Tokens is essential to cope with the diversity and complexities in a scene. We evaluate our density- and scale-aware attention weighting scheme which is designed to augment the scaled-dot-product self-attention and allow the model to dynamically adjust its focus based on the scale of objects and their surrounding density. Findings in Tab. 7 indicate that masking the influence of overlapping RoIs plays a crucial role. This observation aligns with our initial understanding that indiscriminately emphasizing neighboring RoIs, without considering overlap, could lead to skewed attention distributions and potentially impair the model’s ability to accurately discern between distinct objects.
Module mAP50
Pre cls supervision w/o 69.86